OpenStudy (anonymous):

Absolute value 7-4x over 7 absolute value less than or equal to 3

myininaya (myininaya):

\[|\frac{7-4x}{7}|\le3\] ?

myininaya (myininaya):

is that right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[|7-4x|\le21\] \[7-4x \le 21\] , \[7-4x \ge -21\] \[-4x \le 21-7 \] , \[ -4x \ge -21-7\] \[-4x \le 14\] , \[-4x \ge -28\] \[x \ge \frac{14}{-4}\] , \[x \le \frac{-28}{-4}\] \[x \ge \frac{-7}{2}\] , \[x \le 7\] *~~~~* ____|_____|_____ -7/2 7 [-7/2,7] i think it is out of order i tried to do in columns but oh well im off to bed hope this helps

OpenStudy (anonymous):

Guys, is Latex working for you?!